OpenStudy (anonymous):

Write the equation of the line parallel to x = -16 and passes through the point (9, −10).

OpenStudy (anonymous):

Do I use y-y1=m(x-x1) equation??

OpenStudy (imstuck):

Yes you use that. It is the point-slope equation. Use the slope they gave you and then the point they gave you to write the equation.

OpenStudy (anonymous):

m would be x=-16 right?

OpenStudy (imstuck):

Wait...you can't have a slope of x = -16. That's an undefined slope. The slope is in the form y = mx + b. y = ...not x =

OpenStudy (imstuck):

An "x = " line is a vertical line. Vertical lines do not have a slope...

OpenStudy (anonymous):

so i would use y=mx+b equation to get m

OpenStudy (imstuck):

no you can't use that. You would have to know the y intercept and you don't. I don't know what to tell you. x = -16 is an undefined slope. There is a problem within that statement either due to inaccuracy or something.

OpenStudy (anonymous):

Here are the multiple choice answers. Help someone. x = -10 y = -10 x = 9 y = -6

OpenStudy (imstuck):

It's x = 9. That is the only line that is parallel to x = -16 and that goes through your point (9, -10).

OpenStudy (imstuck):

In this case you do not use the point-slope form, as you can see. You just have to satisfy the conditions that the line is parallel to another x line, and that it goes through the x coordinate that they gave you.